1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern,

The Werewolf  Der Werwolf  El Hechicero

tradukita de Alexander Gross tradukita de Hans Eiseneck

A Werewolf, troubled by his name, Left wife and brood one night and ca-

me To a hidden graveyard to enlist The aid of a long-dead philologist.

Ein Werwolf eines Nachts ent-

wich von Weib und Kind und sich

begab an eines Dorfschullehrers

Grab und bat ihn: �Bitte, beuge

mich!“ Un hechicero, con esposa vino un día a la fosa de su maestro, y al pro-

fesor pidió: “¿Me cuentas?

¡Por favor!”

“Oh sage, wake up, please don’t bera-

te me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and

then was ready.

Der Dorfschulmeister stieg

hinauf auf seines Blechschilds Mes-

singknauf und sprach zum Wolf, der sei-

ne Pfoten geduldig kreuzte vor dem To-

ten:

El profesor sin vacilar subió y comenzó a hab-

lar. Desde allá de su letrero se dirigió al hechicero:

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present

tense, And ‘Iswolf,’ ‘Arewolf’ in this same

sense.”

�Der Werwolf“ – sprach der

gute Mann, �des Werwolfs, Genitiv so-

dann, dem Werwolf, Dativ, wie

man’s nennt, den Werwolf, - damit hat’s ein

End!“ “El hechiuno” le explica, “el hechidos” después in-

dica, “el hechitres” dice al fin, “con eso basta ¡chi-

quitín!”

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

Dem Werwolf schmeichelten

die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge doch zur Einzahl auch die Mehrzahl

noch!“ El hechicero, conmovido por lo que capta su oído, al maestro ruega, encan-

tado, que le adjunte un quebra-

do.

…  …  …

2/2

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Ut-

terly wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Der Dorfschullehrer aber mus-

ste gestehn, dass er von ihr nichts

wusste. Zwar Wölfe gab’s in großer

Schar, doch �Wer“ gab’s nur im Sinu-

lar.

Pero el profesor confiesa que nada sabe con certe-

za: “Hay muchos hechos. He-

chiceros únicamente hay enteros.”

The Werewolf knew better–his sons

still slept At home, and homewards now he

crept, Happy, humble, without apology For such folly of philology.

Der Wolf erhob sich

tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehrter

eben, so schied er dankend und er-

geben.

El brujo se volvió muy

triste: “Mi cara mitad, ¿no la vi-

ste?” Ya que no era sabio con gratitud se despidió.

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en hispana de Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .